OpenStudy (anonymous):

4x=y^2 this question is racking my brain. what steps r needed to solving it?

OpenStudy (anonymous):

I m solving for x

OpenStudy (anonymous):

divide by 4

OpenStudy (anonymous):

on both sides? how can u divide y squared by 4?

OpenStudy (turingtest):

you can divide any variable by any number except zero. simply write it as a fraction as you would with two numbers; variable in the numerator, number in the denominator.

OpenStudy (turingtest):

i.e. 12m=z^3 m=z^3/12

OpenStudy (turingtest):

oh I'm sorry, I didn't realize this was your question XD yeah you had the right answer